Suppose the risk-free rate is 3.61% and an analyst assumes a market risk premium of 5.35%. Firm A just paid a dividend of $1.40 per share. The analyst estimates the β of Firm A to be 1.48 and estimates the dividend growth rate to be 4.44% forever. Firm A has 263.00 million shares outstanding. Firm B just paid a dividend of $1.66 per share. The analyst estimates the β of Firm B to be 0.76 and believes that dividends will grow at 2.31% forever. Firm B has 185.00 million shares outstanding. What is the value of Firm B?

Answer format: Currency: Round to: 2 decimal places.

Suppose the risk-free rate is 2.22% and an analyst assumes a market risk premium of 7.39%. Firm A just paid a dividend of $1.00 per share. The analyst estimates the β of Firm A to be 1.35 and estimates the dividend growth rate to be 4.70% forever. Firm A has 266.00 million shares outstanding. Firm B just paid a dividend of $1.95 per share. The analyst estimates the β of Firm B to be 0.86 and believes that dividends will grow at 2.38% forever. Firm B has 193.00 million shares outstanding. What is the value of Firm A?

Answer format: Currency: Round to: 2 decimal places.

Caspian Sea Drinks needs to raise $56.00 million by issuing additional shares of stock. If the market estimates CSD will pay a dividend of $1.48 next year, which will grow at 3.53% forever and the cost of equity to be 11.41%, then how many shares of stock must CSD sell?

Answer format: Number: Round to: 0 decimal places.

i would really appreciate the help!! :)
Answer #1

Ans.1). Cost of equity (ke) = risk-free rate + (beta*market risk premium)

= 3.61% +(0.76*5.35%) = 7.68%

Current price (P) = D0*(1+g)/(ke -g) where D0 = dividend = 1.66 and g = growth rate = 2.31%

P = 1.66*(1+2.31%)/(7.68%-2.31%) = 31.65

Firm B value = P*number of shares = 31.65*185,000,000 = $5,855,274,133.43 (Note: This can hold true only if we assume that the firm has no debt as there is no information given in the question about debt.)

Ans.2). Cost of equity (ke) = risk-free rate + (beta*market risk premium)

= 2.22%+(1.35*7.39%) = 12.20%

Current price (P) = D0*(1+g)/(ke -g) where D0 = dividend = 1.00 and g = growth rate = 4.70%

P = 1.00*(1+4.71%)/(12.20%-4.70%) = 13.97

Firm A value = P*number of shares = 13.97*266,000,000 = $3,715,093,710.40 (Again, the assumption is that the firm is unlevered.)

Ans.3). Price per share (P) = D1/(ke-g) = 1.48/(11.41%-3.53%) = 18.78

Number of shares that must be sold = capital needed/price per share = 56,000,000/18.78 = 2,981,621.62 or 2,981,622 (rounded to the nearest whole number)